(2a+3b)(a–b) = 2a^2 –3b^2 ?????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (2a+3b)(a–b) = 2a(a–b)+3b(a–b)
= 2a^2–2ab+3ab–3b^2
= 2a^2+ab–3b^2
≠2a^2–3b^2 = RHS
The correct statement is (2a +3b)(a–b) = 2a^2+ab–3b^2

Thanks